# Generate, Transduct, Adapt: Iterative Transduction with VLMs

### v1

@@ 4 Experiments
To generate attributes, we use Llama-3.1 with a maximum token length of 1000. Initial experiments on the CUB dataset showed similar performance to GPT-4, so we selected Llama due to its open-source availability. For each pair of classes, we use the prompt described in § 3.2 to generate attributes. The threshold $\alpha$ for selecting the confused images is set to 0.1, and the hyperparameter $\beta$ is adjusted so that the cumulative count $\mathcal{CC}$ includes 5% of the most confused images. We run GTA-CLIP for 30 iterations (i.e., $T=30$ in Algorithm 1) and select the top $k=8$ images per class for fine-tuning using the labels in $\mathbf{z}$. These parameters remain fixed across all datasets, and we found our approach robust to these choices within a reasonable range (see Appendix for a sensitivity analysis). For our experiments, we use the ViT-B/32, ViT-B/16, and ViT-L/14 architectures of CLIP. Fine-tuning is performed with the AdamW optimizer, using betas of $(0.9,0.98)$, an epsilon of 1E-6, and a batch size of 32. We set a learning rate of $\gamma=$ 2E-7 and weight decay of $\lambda=$ 1E-4 for the Transformer layers of the image and text encoders, and $\gamma=$ 1E-6 and $\lambda=$ 1E-4 for the final linear projection layers. All results are reported in terms of Top-1 accuracy, averaged over 3 runs.
Resource Requirements.
All experiments are run on a node with a single A100 GPU and a batch size of 32. Our fine-tuning process is highly efficient as it uses only 8 examples per class. While generating comparative attributes with LLama is constrained by strict thresholds to manage complexity, it remains the most computationally intensive step. The number of pairs for attribute generation decreases with each epoch, typically reaching zero for most datasets after about 10-15 epochs due to fixed thresholds $\alpha$ and $\beta$. Overall, the total time complexity scales with the number of classes; for example, on the 200-class CUB dataset, GTA-CLIP completes in roughly one hour. Notably, if attribute generation is excluded, the entire setup runs in just 10 mins.
@@ 5 Results
We present results for the zero-shot (§ 5.1), few-shot (§ 5.2), and zero-shot with seen classes (§ 5.3) setting on benchmark datasets, followed by ablation studies (§ 5.4) and a detailed analysis of our method (§ 5.5).
@@ 5.1 Zero-Shot Performance
Table 1 shows the zero-shot performance of GTA-CLIP compared to the CLIP [36] and the current state-of-the-art, TransCLIP [61]. We report accuracy across 12 datasets using different CLIP architecture—ViT-B/32, ViT-B/16, and ViT-L/14—along with the overall average accuracy. GTA-CLIP improves over TransCLIP by 3.96%, 4.01%, and 3.22%and over CLIP by 8.58%, 9.46%, and 7.78% on average using B/32, B/16, and L/14 respectively.
@@ 5.2 Few-Shot Performance
Table 2 shows the few-shot performance of our approach compared to TransCLIP. We report results using the ViT-B/16 architecture with $1$-shot, $4$-shot, and $16$-shot settings denoting the number of labeled training examples per category are provided. We use the TransCLIP-FS reported in their paper. Both ours and TransCLIP can incorporate labeled examples by simply setting the corresponding entries in $\mathbf{z}$ to the one hot vector corresponding to their labels, as described in § 3. Performance is reported on the same set of images in the zero-shot setting.
Like the zero-shot case for CLIP ViT-B/16, we find that in every setting and every choice of $k$-shot, GTA-CLIP outperforms TransCLIP. We find an increase of 3.41%, 3.86%, and 3.01% for $1$-shot, $4$-shot, and $16$-shot, respectively. Once again our method improves performance in every setting, with the most pronounced performance increase using $4$-shot. Trends of improvements align with the zero-shot setting. Interestingly, we find that zero-shot GTA-CLIP outperforms $1$-shot TransCLIP, saving human effort, as labeling even a single example per category can be labor-intensive for certain datasets. Furthermore, we find that the gains from transduction, attribute-guided transduction with adaptation (our approach) complement labeling efforts. This flexibility is of practical value, offering end users multiple ways to improve performance on a target dataset.
@@ 5.3 Zero-Shot Performance with Seen Classes
Previous work has also evaluated zero-shot learning in a setting where labeled data from a related but different set of classes is available during training, with performance evaluated on images from unseen classes. Approaches such as CoCoOp [66], AdaptCLIPZS [40] and CLIP-A-self [24] report results by splitting a dataset’s categories in half, treating the first half as “seen” classes to adapt their model and measuring performance on the “unseen” second half. The results are shown in Table 3. Performance tend to be higher in this setting in comparison to the zero-shot and few-shot experiments, as only half of the classes are considered. [^2] Only the CUB dataset has lower performance as the test split is harder than the overall dataset.
For a straightforward comparison, we initialize the CLIP model with the pre-trained weights from AdaptCLIPZS ViT-B/16 and report the accuracies of TransCLIP and GTA-CLIP on the “unseen” classes of each dataset, using the same framework as the zero-shot setting. Note that this setup does not include a transductive term between training and testing images, thus representing a lower bound on achievable performance. Despite this, we find that GTA-CLIP outperforms prior methods across all five datasets considered: CUB200 [47], Stanford Cars [17], FGVC Aircraft [23], Flowers102 [31], and Food101 [5].
While AdaptCLIPZS, CLIP-A-self, CoOp, CoCoOp, and CLIP-A use an inductive setup, TransCLIP and GTA-CLIP adopt a transductive approach that benefits from having test images available in advance. This results in improvements in similar vein as the zero-shot setting. However, even with domain-specific fine-tuning of CLIP with labels, transductive inference proves advantageous, and our attribute-guided approach yields further improvements–an encouraging result. The improvements over CLIP are substantial, though this setup requires more supervision than the previous settings.
Table 2: Few-shot Results. Performance ($1$-shot, $4$-shot, and $16$-shot) of GTA-CLIP and TransCLIP-FS across datasets using CLIP ViT-B/16 network. We find that GTA-CLIP outperforms TransCLIP-FS in all cases.
@@
Table 3: Zero-shot Results with Seen Classes. In this setting examples from a “seen” classes are used to adapt CLIP ViT-B/16 and evaluated on “unseen” classes. We show the performance of both inductive and transductive approaches. Some techniques such as CLIP-A-self [24] use 3:1 split as opposed to the 1:1 used by other methods on CUB so we do not include their numbers. Tranductive inference remains beneficial, and GTA-CLIP improves over TransCLIP.
@@ 5.4 Ablation Studies
We next aim to quantify the performance contributions of each component in Algorithm 1. In Table 4, we selectively disable components of our method and report the average performance across CUB200 [47], Stanford Cars [17], FGVC Aircraft [23], Flowers102 [31], and Food101 [5]. We also present the relative percentage improvement over CLIP using the CLIP ViT-B/16 architecture in the full test set zero-shot setting. Note that Adapt cannot be applied in isolation in this setting due to its requirement for labeled data (either weak or ground-truth), so it is only included when Transduct and the necessary sample selection strategy are also present.
@@
The Appendix analyzes the effects of more implementation choices of our method. We also include a sensitivity analysis, where we provide performance over the hyperparameters $k,T,\alpha,\beta$ of our method given in Algorithm 1.
5.5 Analysis
@@
Table 5 also visualizes the progression of confusion counts for the top confused class pair (Western Gull, California Gull). The number of images with a top-2 probability difference below $\alpha=0.1$ generally decreases over epochs. The most significant drop occurs between the first and second epochs, highlighting the impact of the newly generated attributes.
Figure 2: t-SNE Plots of Class Attributes. For each category the prototype, initial set of attributes, and the final set of attributes are shown in green, blue, and red respectively. Habitat, relative characteristics, and other distinguishing features are often identified through pairwise comparisons, while the initial attributes tend to describe the prominent visual features. These plots were obtained by mapping the CLIP text embeddings of the attributes using t-SNE. Please zoom in for details and see Appendix for a detailed figure.
@@ 6 Limitations
There are two main limitations to our work. The first is the use of LLMs to generate fine-grained attributes. LLMs are known to sometimes hallucinate data, and there is a risk that some attributes may be incorrect. In our initial experiments, we also found that results vary depending on the LLM used. Additionally, there is an interaction with the benchmark to which the LLM is applied; a higher-performing language model might generate lower-quality fine-grained attributes on a specific benchmark due to a pre-training data mismatch, which can lead to poorer performance with our method. The second limitation is the applicability of transductive learning. This approach is constrained by the extent of knowledge available about the query set before evaluation. In this work, we have not demonstrated robustness to imperfect knowledge of the classes and data distribution. For instance, we implicitly assume that images within a class cluster together, allowing us to model each class as a Gaussian distribution, which may not always be valid.
@@ 7 Conclusion
Despite its limitations, the transductive setting offers a compelling approach for practitioners and domain experts who need precise answers for specific datasets. For instance, an ecologist might be interested in estimating species counts from data gathered via a network of camera traps, while a scientist might want to determine land-cover distribution using satellite imagery. VLMs enable straightforward labeling through language-based descriptions of categories, but their initial accuracy is often insufficient. Our work demonstrates that expanding categories based on attributes, when combined with transductive learning, enables model fine-tuning to achieve significant accuracy improvements. Additionally, this approach offers complementary advantages to traditional labeling methods, such as providing a few labeled examples per class. While we use large language models for convenience, this iterative procedure is naturally suited to human-in-the-loop approaches, allowing practitioners to incrementally add attributes and labels for ambiguous classes. These findings are practically valuable, as they offer end-users multiple pathways to improve labeling precision on their target dataset without investing significant efforts on training dataset curation and model training.
Acknowledgments
Experiments were performed on the University of Massachusetts GPU cluster funded by the Mass. Technology Collaborative. OS and SM were supported in part by a NSF grant #2329927.
@@ 8 Additional Ablations
Table 6: Per-dataset results of Ablation Study. For five datasets on the ViT-B/16 architecture, we present the effect of various components of GTA-CLIP. We use the same conventions as Tab. 4.
@@
Tab. 6 breaks down the Top-1 accuracies across reported in Tab. 4 of the main papers across individual datasets namely CUB, Stanford Cars, FGVC Aircraft, Flowers102, and Food101 datasets using the ViT-B/16 architecture. We observe similar trends for each dataset for all the ablations considered.
@@
MetaCLIP introduces better CLIP architectures by curating training data and scaling training. We switch the base VLM from CLIP to MetaCLIP to take advantage of this and test the generalization of our approach to new architectures. Tab. 7 presents the accuracies for the inductive version of MetaCLIP, TransMetaCLIP (the TransCLIP method applied to MetaCLIP), and GTA-MetaCLIP (our method applied to MetaCLIP). The experiments are conducted using the ViT-B/16 architecture of MetaCLIP across CUB, Stanford Cars, FGVC Aircraft, Flowers102, and Food101 datasets. We observe consistent improvements in the case of MetaCLIP too. On average, over the five datasets, we see an improvement of 6.8% over MetaCLIP, and an improvement of 2.7% over TransMetaCLIP on using our method. This is similar to our improvements of 7.0% and 3.3% on the corresponding baselines with CLIP.
@@
For all the results in the main paper, we used Llama-3.1 as the LLM model for dynamic attribute generation. Now we explore using GPT4o as the LLM in Tab. 8. We observe that the accuracy remains similar on average over the CUB, Stanford Cars, FGVC Aircraft, Flowers102, and Food101 datasets on ViT-B/16. Thus, using Llama-3.1 is a more cost-effective choice for dynamic attribute generation due to its open-source nature.
Removing the internal call to Transduct in GenerateAttributes. We remove the internal transductive inference update call (see § 3.2) in GenerateAttributes and present the results over five datasets using the ViT-B/16 architecture in Tab. 9. We observe that on average the accuracy drops on removing this call to Transduct. For Aircraft, we observe that the accuracy slightly improves on dropping this Transduct call, however for Cars we see a significant decrease.
@@ 9 Sensitivity Analysis
Similarly, in Tab. 11 we show the performance of GTA-CLIP when varying the probability threshold for determining confusing pairs of classes, $\alpha$. For the whole experiment, we fix $k=8,T=30$ and sweep over $alpha$. We find that each benchmark has it’s own ideal $\alpha$ value, namely that no two benchmark’s max performances share an common $alpha$. Surprisingly, we see that $\alpha=17.5\%$, which does not perform the best on any benchmark, has the highest average value. We also conclude that GTA-CLIP has a greater insensitivity to the choice of $\alpha$ as compared to $T$ but similar to $k$. Namely we find that the spread of $\alpha$ to be $67.61-67.06=0.55$, $T$ to be $67.46-65.89=1.57$, and $k$ to be $67.20-66.76=0.44$. Finally, we find that the minimum performance increase by introducing GenerateAttributes is at $\alpha=5.0\%$ with a gain of $67.10-66.76=0.34$. In other words, adding any amount of comparative attribute generation improves performance.

### v2

@@ 4 Experiments
To generate attributes, we use Llama-3.1 with a maximum token length of 500. All experiments are run on a single A100 GPU. For each pair of classes, we use the prompt described in § 3.2 to generate attributes. The threshold $\alpha$ for selecting the confused images is set to 0.1, and the hyperparameter $\beta$ is adjusted so that the cumulative count $\mathcal{CC}$ includes 5% of the most confused images. We run GTA-CLIP for 30 iterations (i.e., $T=30$ in Algorithm 1) and select the top $k=8$ images per class for fine-tuning using the labels in $\mathbf{z}$. These parameters remain fixed across all datasets, and we found our approach robust to these choices within a reasonable range (see Appendix for a sensitivity analysis).
For our experiments, we use the ViT-B/32, ViT-B/16, and ViT-L/14 architectures of OpenAI’s CLIP models (Performance using CLIP models from Meta are included in the Appendix). Fine-tuning is performed with the AdamW optimizer, using betas of $(0.9,0.98)$, an epsilon of 1E-6, and a batch size of 32. We set a learning rate of $\gamma=$ 2E-7 and weight decay of $\lambda=$ 1E-4 for the Transformer layers of the image and text encoders, and $\gamma=$ 1E-6 and $\lambda=$ 1E-4 for the final linear projection layers. All results are reported in terms of Top-1 accuracy, averaged over 3 runs.
@@ 5 Results
We present results for zero-shot (§ 5.1), few-shot (§ 5.2), and zero-shot with seen classes (§ 5.3) setting on various datasets, followed by ablation studies (§ 5.4) and a detailed analysis of our method (§ 5.5). We also provide further experiments in our Appendix, including similar performance improvements using MetaCLIP (§ 9), sensitivity analysis (§ 10), to show robustness of our method, and detailed visualization of generated attributes (§ 11).
@@ 5.1 Zero-Shot Performance
Table 1 shows the zero-shot performance of GTA-CLIP compared to the CLIP [36] and the current state-of-the-art, TransCLIP [61]. We report accuracy across 12 datasets using different CLIP architecture—ViT-B/32, ViT-B/16, and ViT-L/14—along with the overall average accuracy. GTA-CLIP improves over TransCLIP by 3.96%, 4.01%, and 3.22%% and over CLIP by 8.58%, 9.46%, and 7.78% on average using B/32, B/16, and L/14 respectively.
@@ 5.2 Few-Shot Performance
Table 2 shows the few-shot performance of our approach compared to TransCLIP. We report results using the ViT-B/16 architecture with $1$-shot, 44-shot, and $16$-shot settings denoting the number of labeled examples per class. We use the TransCLIP-FS [61] setting for this. Both ours and TransCLIP can incorporate labeled examples by simply setting the corresponding entries in $\mathbf{z}$ to the one hot vector corresponding to their labels, as described in § 3. Performance is reported on the same set of images in the zero-shot setting.
Like the zero-shot case for CLIP ViT-B/16, we find that in every setting and every choice of $k$-shot, GTA-CLIP outperforms TransCLIP. We find an increase of 3.41%, 3.86%, and 3.01% for $1$-shot, 44-shot, and $16$-shot, respectively. We observe the most pronounced performance increase for $4$-shot. Trends of improvements align with the zero-shot setting. Interestingly, we find that zero-shot GTA-CLIP outperforms $1$-shot TransCLIP, saving human effort, as labeling even a single example per category can be labor-intensive for certain datasets. Furthermore, we find that the gains from transduction, attribute-guided transduction with adaptation (our approach) complement labeling efforts. This flexibility is of practical value, offering end users multiple ways to improve performance on a target dataset.
@@ 5.3 Zero-Shot Performance with Seen Classes
Previous work has also evaluated zero-shot learning in a setting where labeled data from a related but different set of classes is available during training, while performance is evaluated on images from unseen classes. Approaches such as CoCoOp [66], AdaptCLIPZS [40] and VDT [24] report results by splitting a dataset’s categories in half, treating the first half as “seen” classes to adapt their model and measuring performance on the “unseen” second half. The results are shown in Table 3. Performance tends to be higher in this setting in comparison to the zero-shot and few-shot experiments, as only half of the classes are considered [^2] Only the CUB dataset has lower performance as the test split is harder than the overall dataset. and a smaller domain shift.
For a straightforward comparison, we initialize the CLIP model with the pre-trained weights from AdaptCLIPZS and report the accuracies of TransCLIP and GTA-CLIP on the “unseen” classes of each dataset, using the same framework as the zero-shot setting. Note that this setup does not include a transductive term between training and testing images, thus representing a lower bound on achievable performance. Despite this, we find that GTA-CLIP outperforms prior methods across all five datasets considered: CUB [47], Stanford Cars [17], FGVC Aircraft [23], Flowers102 [31], and Food101 [5].
While AdaptCLIPZS, VDT, CoOp, CoCoOp, and CLIP-A use an inductive setup, TransCLIP and GTA-CLIP adopt a transductive approach that benefits from having test images available in advance. This results in improvements in similar vein as the zero-shot setting. However, even with domain-specific fine-tuning of CLIP with labels, transductive inference proves advantageous, and our attribute-guided approach yields further improvements–an encouraging result. The improvements over CLIP are substantial, though this setup requires more supervision than the previous settings.
Table 2: Few-shot Results. Performance (11-shot, 44-shot, and $16$-shot) of GTA-CLIP and TransCLIP-FS across datasets using CLIP ViT-B/16 network. We find that GTA-CLIP outperforms TransCLIP-FS in all cases.
@@
Table 3: Zero-shot Results with Seen Classes. In this setting examples from a “seen” classes are used to adapt CLIP ViT-B/16 and evaluated on “unseen” classes. We compare both inductive and transductive approaches. Some techniques such as VDT [24] use 3:1 split as opposed to the 1:1 used by other methods on CUB so we do not include their numbers. Tranductive inference remains beneficial, and GTA-CLIP improves over TransCLIP.
@@ 5.4 Ablation Studies
We next aim to quantify the performance contributions of each component in Algorithm 1. In Table 4, we selectively disable components of our method and report the average performance over five datasets.
@@
5.5 Class Confusion and Attribute Space
@@
Table 4: Ablation Study. Ablation study of the components of GTA-CLIP using ViT-B/16. Average Top-1 accuracy across five datasets is shown (see Appendix for the full table). Attributes $\mathcal{A}=\{$$\text{S},\text{D}\}$ refer to no, static, and dynamic attributes, respectively. No attributes corresponds to standard CLIP, while static and dynamic refer to the initial set of attributes and confusion-driven attributes respectively. The first row shows the performance of CLIP, and the third row shows the performance of TransCLIP. Simply generating attributes leads to insignificant improvement in performance on these fine-grained datasets (row 2), but it improves transductive inference and subsequent adaptation. Dynamic attribute generation provides additional benefits (last row).
Attributes Transduct Adapt Acc. $\Delta$ CLIP
@@
Table 5 also visualizes the progression of confusion counts for the top confused pair (Western Gull, California Gull). The number of images with a probability difference below $\alpha=0.1$ generally decreases over epochs. The most significant drop occurs between the first and second epochs, highlighting the impact of the newly generated attributes.
@@
5.6 Computational Cost
For attribute expansion, we explore both open-source Llama-3.1-8b and GPT-4o, and find their effect on performance to be similar (see Table 8 in Appendix). The number of class pairs requiring attribute expansion decreases with each epoch, reaching zero for most datasets after about 10 epochs due to the chosen thresholds $\alpha$ and $\beta$, ensuring that attributes are not regenerated for duplicate class pairs.
In the CUB dataset, a total of approximately 30 pairs of confusing classes were selected over 30 iterations for prompting the LLM. Across datasets of various sizes, the number of sampled class pairs remains within a similar range, with the Flowers dataset requiring attribute expansion for only three pairs. For larger datasets such as ImageNet, we lower $\alpha$ to 0.05 to keep attribute expansion computationally feasible. Running Llama-3.1-8b on our hardware (a single A100) takes less than 10 minutes for all selected pairs, while using GPT-4o via API calls takes under 2 minutes and costs less than \$1.
Our fine-tuning process is highly efficient, using only 8 examples per category. For a dataset like CUB, which contains approximately 12k images across 200 classes, 30 iterations of fine-tuning and transduction take less than 10 minutes on a single A100 GPU.
Overall, the runtime cost of GTA-CLIP is approximately 10-20 minutes higher than that of CLIP on most datasets. However, these additional costs can be justified given the performance gains, since manually labeling even a small fraction of the dataset would require significantly more time, such as in fine-grained domains.
Figure 2: t-SNE Plots of Class Attributes. For each category the prototype, initial set of attributes, and the final set of attributes are shown in green, blue, and red respectively. Habitat, relative characteristics, and other distinguishing features are often identified through pairwise comparisons, while the initial attributes tend to describe the prominent visual features. These plots were obtained by mapping the CLIP text embeddings of the attributes using t-SNE. Please see the Appendix for detailed figures.
@@ 6 Limitations
There are two main limitations to our work. The first is the use of LLMs to generate fine-grained attributes. LLMs are known to hallucinate data, posing a risk of generating incorrect attributes. However, similar to AdaptCLIPZS [40], which demonstrated through human evaluation that the generated attributes are highly accurate, we did not find this to be a significant issue for the datasets we considered. Additionally, this step can be guided by domain experts.
The second limitation is the applicability of the transductive setup. It requires access to the entire test set at once, which may not always be feasible (e.g., in a streaming setting). Additionally, we have not demonstrated robustness to imperfect knowledge of class distributions and data structure. For instance, we assume that images within a class cluster together, allowing us to model each class as a Gaussian distribution, which may not always be true.
@@ 7 Conclusion
The transductive setting offers a compelling approach for practitioners and domain experts who need precise answers for specific datasets. For instance, an ecologist might be interested in estimating species counts from data gathered via a network of camera traps, while a scientist might want to determine land-cover distribution using satellite imagery. VLMs enable straightforward labeling through language-based descriptions of categories, but their initial accuracy is often insufficient. Our work demonstrates that expanding categories based on attributes, when combined with transductive learning, enables model fine-tuning to achieve significant accuracy improvements. Additionally, this approach offers complementary advantages to traditional labeling methods, such as providing a few labeled examples per class. While we use large language models for convenience, this iterative procedure is naturally suited to human-in-the-loop approaches, allowing practitioners to incrementally add attributes and labels for ambiguous classes. These findings are practically valuable, as they offer end-users multiple pathways to improve labeling precision on their target dataset without investing significant efforts on training dataset curation and model training.
8 Acknowledgements
The research is supported in part by grant #2329927 from the National Science Foundation (USA). Our experiments were performed on the GPU cluster funded by the Mass. Technology Collaborative.
@@ 9 Additional Ablations
Table 6: Per-dataset results of Ablation Study. For five datasets on the ViT-B/16 architecture, we present the effect of various components of GTA-CLIP. We use the same conventions as Table 4.
@@
Table 6 breaks down the Top-1 accuracies across reported in Table 4 of the main papers across individual datasets namely CUB, Stanford Cars, FGVC Aircraft, Flowers102, and Food101 datasets using the ViT-B/16 architecture. We observe similar trends for each dataset for all the ablations considered.
@@
MetaCLIP introduces better CLIP architectures by curating training data and scaling training. We switch the base VLM from CLIP to MetaCLIP to take advantage of this and test the generalization of our approach to new architectures. Table 7 presents the accuracies for the inductive version of MetaCLIP, TransMetaCLIP (the TransCLIP method applied to MetaCLIP), and GTA-MetaCLIP (our method applied to MetaCLIP). The experiments are conducted using the ViT-B/16 architecture of MetaCLIP across CUB, Stanford Cars, FGVC Aircraft, Flowers102, and Food101 datasets. We observe consistent improvements in the case of MetaCLIP too. On average, over the five datasets, we see an improvement of 6.8% over MetaCLIP, and an improvement of 2.7% over TransMetaCLIP on using our method. This is similar to our improvements of 7.0% and 3.3% on the corresponding baselines with CLIP.
@@
For all the results in the main paper, we used Llama-3.1 as the LLM model for dynamic attribute generation. Now we explore using GPT4o as the LLM in Table 8. We observe that the accuracy remains similar on average over the CUB, Stanford Cars, FGVC Aircraft, Flowers102, and Food101 datasets on ViT-B/16. Thus, using Llama-3.1 is a more cost-effective choice for dynamic attribute generation due to its open-source nature.
Removing the internal call to Transduct in GenerateAttributes. We remove the internal transductive inference update call (see § 3.2) in GenerateAttributes and present the results over five datasets using the ViT-B/16 architecture in Table 9. We observe that on average the accuracy drops on removing this call to Transduct. For Aircraft, we observe that the accuracy slightly improves on dropping this Transduct call, however for Cars we see a significant decrease.
@@ 10 Sensitivity Analysis
Similarly, in Tab. 11 we show the performance of GTA-CLIP when varying the probability threshold for determining confusing pairs of classes, $\alpha$. For the whole experiment, we fix $k=8,T=30$ and sweep over alphaalpha. We find that each benchmark has it’s own ideal $\alpha$ value, namely that no two benchmark’s max performances share an common alphaalpha. Surprisingly, we see that $\alpha=17.5\%$, which does not perform the best on any benchmark, has the highest average value. We also conclude that GTA-CLIP has a greater insensitivity to the choice of $\alpha$ as compared to $T$ but similar to $k$. Namely we find that the spread of $\alpha$ to be 67.6167.06=0.5567.61-67.06=0.55, $T$ to be 67.4665.89=1.5767.46-65.89=1.57, and $k$ to be 67.2066.76=0.4467.20-66.76=0.44. Finally, we find that the minimum performance increase by introducing GenerateAttributes is at $\alpha=5.0\%$ with a gain of 67.1066.76=0.3467.10-66.76=0.34. In other words, adding any amount of comparative attribute generation improves performance.